Write an equation for an ellipse if the endpoints of the major axis are at (1,6) and (1,-6) and the endpoints of the minor axis are at (5,0) and (-3,0).
I'm lost
thanks

Answer by venugopalramana(3286)

Write an equation for an ellipse if the endpoints of the major axis are at (1,6) and (1,-6)
HENCE CENTRE IS (1+1)/2,(6-6)/2=(1,0)...H=1...K=0
MAJOR AXIS LENGTH = 6+6=12=2A....OR.....A=6
and the endpoints of the minor axis are at (5,0) and (-3,0)
MINOR AXIS LENGTH =5+3=8=2B.....OR.....B=3
HENCE EQN.OF ELLIPSE IS
(X-H)^2/A^2 + (Y-K)^2/B^2 =1
(X-1)^2/36 + Y^2/9 =1
I'm lost
thanks